# Stabilizer Entanglement Distillation and Efficient Fault-Tolerant Encoder

### v1

@@ VII Conclusion
In this paper, we first review stabilizer entanglement distillation and apply it to the well-known Bennett’s and Deutsch’s $2$-to-$1$ distillation protocols, mapping these protocols onto a $\left[2,1,2\right]$ classical error detection code. We show that bilateral Hadamard gates on the distillation output can alternate Pauli-$X$ and Pauli-$Z$ errors while preserving the standard Bell states. Therefore, two sequential applications of these protocols can effectively address any single-qubit errors. Building on this insight, we extend the $2$-to-$1$ distillation protocol to higher-rate $n$-to-$\left(n-1\right)$ protocols, which achieve higher yield for high-fidelity input Bell states. We also present a method to adapt any classical error correcting code to an entanglement distillation protocol using the two-iteration process. The first iteration corrects bit-flip errors, then Hadamard gates convert phase-flip errors to bit-flip errors, which the second iteration corrects. The errors in both iterations are classical, so the error-correcting capability matches that of the classical code used.
Furthermore, we propose a constant-depth decoder for stabilizer codes. We apply it to a stabilizer entanglement distillation protocol based on quantum convolutional codes, demonstrating reduced circuit complexity compared to protocols using a unitary decoding circuit. We also develop a fault-tolerant encoding and decoding protocol for qLDPC codes and surface codes using constant-depth circuits. This protocol combines techniques from stabilizer entanglement distillation, constant-depth decoder, and quantum teleportation. These operations are feasible with state-of-the-art reconfigurable atom arrays, and the encoder surpasses the conventional depth limits for surface codes, which are lower bounded by $\mathcal{O}\left(\log\left(n\right)\right)$.
Our work directly applies to entanglement routing protocols in quantum networks [115, 116]. The $n$-to-$\left(n-1\right)$ recurrence protocol integrates seamlessly with current entanglement distillation technologies. Measurement-based decoding reduces circuit depth and minimizes latency, thereby enhancing the performance of quantum repeaters. Overall, our work integrates stabilizer formalism, measurement-based quantum computing, and entanglement distillation, advancing both quantum communication and computing, and contributing to the development of practical and efficient quantum information processing techniques.
@@ Appendix A Bell-State Matrix Identity
We show the Bell-state matrix identity for $n$ Bell states. The standard Bell state is defined by
@@
Appendix B Integrating Stabilizer Measurements, Error Correction, and Decoding
@@
We can integrate stabilizer measurements with single-qubit measurement decoding. Consider a stabilizer code in its standard form. The $i$th generator, for $i<r$, consists of a single bit flip on the $i$th qubit. After measuring $g_{i}$, the following stabilizer and single-qubit measurements use only Pauli-$Z$s on the $i$th qubit. This allows us to replace $Z_{i}$ in these stabilizers with its measurement outcome, reducing the number of two-qubit gates.
For example, consider the $\left[\!\left[5,1,3\right]\!\right]$ code. Its stabilizer generators are
We can integrate stabilizer measurements and single-qubit measurement decoding into a circuit as shown in Figure 7. This circuit yields both stabilizer parities and decoding parities. Using this information, Bob can determine the recovery operators for his last $k$ qubits, resulting in distilled Bell states.
Appendix C Decoding Quantum Convolutional Codes with Laurent Series
@@
$X_{L}^{i}$ $=$ $\bigotimes_{l:q_{l}=1}^{m}X_{\left(l+1\right)n-k+i}\;,$
$Z_{L}^{i}$ $=$ $\bigotimes_{l:q_{l}=1}^{m}Z_{\left(l+1\right)n-k+i}\;.$
These operators require further decoding. However, it can be shown that the logical qubit originates from a catastrophic encoding circuit. For example, with $Q\left(D\right)=1+D+D^{2}$, the logical qubits are encoded by the following circuit
@@
$\begin{matrix}\left(-1\right)^{a_{x}^{1}+a_{x}^{2}+b_{x}^{1}+b_{x}^{2}}X_{3}\otimes X_{3}\;,&\left(-1\right)^{a_{z}^{1}+a_{z}^{2}+b_{z}^{1}+b_{z}^{2}}Z_{3}\otimes Z_{3}\\ \left(-1\right)^{a_{x}^{2}+b_{x}^{2}}X_{3}X_{4}\otimes X_{3}X_{4}\;,&\left(-1\right)^{a_{z}^{2}+b_{z}^{2}}Z_{3}Z_{4}\otimes Z_{3}Z_{4}\\ X_{3}X_{4}X_{5}\otimes X_{3}X_{4}X_{5}\;,&Z_{3}Z_{4}Z_{5}\otimes Z_{3}Z_{4}Z_{5}\\ \vdots&\vdots\end{matrix}\;.$
Here, $a_{x}^{1}$ and $b_{x}^{1}$ are the measurement outcomes of $X_{-1}X_{1}$ by Alice and Bob, respectively. $a_{x}^{2}$ and $b_{x}^{2}$ are the outcomes of $X_{0}X_{2}$, $a_{z}^{1}$ and $b_{z}^{1}$ are the outcomes of $Z_{-1}Z_{1}$, and $a_{z}^{2}$ and $b_{z}^{2}$ are the outcomes of $Z_{0}Z_{2}$. By recursively calculating the parities for subsequent physical Bell states, these logical Bell states are decoded.
Errors in parity measurements can propagate indefinitely. To address this, we segment the sequence into blocks of $N$ frames. Measuring parities independently for each block confines error propagation to a length of $N$. For a general polynomial $Q\left(D\right)=\sum_{l=0}^{m}q_{l}D^{l}$, the entanglement-assisted approach consumes $m$ pre-shared Bell states and the first $m$ Bell states per block, yielding $N-m$ decoded physical Bell states. The overhead, defined as the ratio of consumed to output Bell states, is $\frac{2m}{N-m}$, approaching zero as $N$ increases.
Appendix D Deriving the Standard Form for the Rate-1/3 Convolutional Code
@@
This process results in the standard form as specified in Equation 17. Note that during Gaussian elimination, multiplication factors representing finite shifts are preferred. Using a Laurent series as a multiplication factor, such as $\frac{1}{1+D}=1+D+D^{2}+\cdots$, can result in a standard form consisting of Laurent series entries, increasing decoding complexity. There is no general algorithm, to our knowledge, that guarantees producing a standard form consisting solely of polynomial entries.
Appendix E Modeling Errors in Measurement-Based Decoding
@@
$X_{L}^{i}$ $=$ $\left(-1\right)^{\alpha_{x}^{i}}X_{n-k+i}\;,$
$Z_{L}^{i}$ $=$ $\left(-1\right)^{\alpha_{z}^{i}}Z_{n-k+i}\;,$
where $\alpha_{x}^{I}=\sum_{j:C_{ij}^{T}=1}a_{j}+\sum_{j:E_{ij}^{T}=1}a_{r+j}$ and $\alpha_{z}^{i}=\sum_{j:\left(A_{2}^{T}\right)_{ij}=1}a_{j}$, as defined in Equations 10 and 11. Noise from single-qubit measurements leads to phase errors. $\alpha_{x}^{i}$ errors yield a $Z$ error, $alpha_{z}^{i}$ errors yield an $X$ error, and errors in both $\alpha_{x}^{i}$ and $alpha_{z}^{i}$ yield a $Y$ error on the decoded qubit.

### v2

@@ VII Conclusion
In this paper, we first review stabilizer entanglement distillation and apply it to the well-known $2$-to-$1$ distillation protocols of Bennett et al. and Deutsch et al., mapping these protocols onto a $\left[2,1,2\right]$ classical error detection code. We show that bilateral Hadamard gates on the distillation output can alternate Pauli-$X$ and Pauli-$Z$ errors while preserving the standard Bell states. Therefore, two sequential applications of these protocols can effectively address any single-qubit errors. Building on this insight, we extend the $2$-to-$1$ distillation protocol to higher-rate $n$-to-$\left(n-1\right)$ protocols, which achieve higher yield for high-fidelity input Bell states. We also present a method to adapt any classical error-correcting code to an entanglement distillation protocol using the two-iteration process. The first iteration corrects bit-flip errors, then Hadamard gates convert phase-flip errors to bit-flip errors, which the second iteration corrects. The errors in both iterations are classical, so the error-correcting capability matches that of the classical code used.
Furthermore, we propose a constant-depth decoder for stabilizer codes and apply it to a stabilizer entanglement distillation protocol based on quantum convolutional codes, achieving reduced circuit complexity compared to unitary decoding. We also develop an efficient fault-tolerant encoding and decoding protocol for surface codes, discussing its potential and challenges in extending to broader QLDPC codes with higher encoding rates. This protocol combines techniques from stabilizer entanglement distillation, constant-depth decoders, and quantum teleportation. These operations are feasible with state-of-the-art reconfigurable atom arrays, and the encoder surpasses the conventional depth limits for surface codes, which are lower bounded by $\Omega\left(\log\left(n\right)\right)$.
Our work directly applies to entanglement routing protocols in quantum networks [134, 135]. The $n$-to-$\left(n-1\right)$ protocol with small values of $n$ could potentially be demonstrated experimentally in the near future. Measurement-based decoding reduces circuit depth and minimizes latency, thereby enhancing the performance of quantum repeaters. Overall, our work integrates stabilizer formalism, measurement-based quantum computing, and entanglement distillation, advancing both quantum communication and computing, and contributing to the development of practical and efficient quantum information processing techniques.
@@ Appendix A Bell-State Matrix Identity
We show the Bell-state matrix identity for $n$ standard Bell states. The standard Bell state is defined by
@@
We demonstrate the equivalence of the bilateral and single-lateral depolarizing channels on the standard Bell state using the Bell-state identity. Starting with the bilateral depolarizing channel, we assume a channel on both side that applies a Pauli operator from $\{X,Y,Z\}$, each with probability $p_{i}=p/3$, or an identity operator with probability $p_{0}=1-p$. The Kraus operators for this channel are given by $\{\sqrt{p_{i}p_{j}}E_{j}\otimes E_{i}\}$, where $E_{0}=I$ and $E_{i}\in\{X,Y,Z\}$ for $i=1,2,3$. Applying these operators to the Bell state $\left|\Psi^{+}\right\rangle$, we use the Bell matrix identity to express them as $\{\sqrt{p_{i}p_{j}}I\otimes E_{i}E_{j}^{T}\}$. For simplicity, we rewrite the equivalent Kraus operators as $\{\sqrt{p_{i}p_{j}}E_{i}E_{j}\}$, omitting the operator on Alice’s side and applying only to Bob’s qubit. Any minus sign from Pauli-$Y$ operator is absorbed when applied as a superoperator. The Kraus operators fall into three cases:
  1. 1.
    $i=j=0$: The Kraus operator is $\{\left(1-p\right)I\}$.
  2. 2.
    Either $i$ or $j=0$: The Kraus operators are $\{\sqrt{\frac{2}{3}p\left(1-p\right)}E_{k}\}$ for $k=1,2,3$.
  3. 3.
    Both $i$ and $j\neq 0$: The Kraus operators are $\{\frac{\sqrt{3}}{3}pI,\frac{\sqrt{2}}{3}pE_{k}\}$ for $k=1,2,3$.
Summing these terms yields a single-lateral depolarizing channel with Kraus operators $\{\sqrt{1-p^{\prime}}I,\sqrt{\frac{p^{\prime}}{3}}E_{i}\}$, where $p^{\prime}=2p-\frac{4}{3}p^{2}$. Thus, the bilateral depolarizing channel is equivalent to a single-lateral depolarizing channel with an adjusted error probability. The single-lateral model, simplifying simulations, was used. Figure 3, as a function of input and output fidelity, automatically captures the adjusted error probability, remaining unaffected by the choice of channel model.
Appendix B Adaptive Distillation with Constant Yield Rate
In this appendix, we demonstrate that iterative applications of the $\left[\!\left[n^{2},\left(n-1\right)^{2},2\right]\!\right]$ distillation protocol, with adaptive block size $n$ based on input fidelity, achieve a constant overall yield rate. We compute the yield rate $R$ for bit-flip error correction. For both bit-flip and phase-flip errors, the yield rate is given by $R^{2}$. The Bell states are characterized by their infidelity (error rate) $p=1-\mathcal{F}$, where $\mathcal{F}$ denotes the fidelity. In the following calculations, we consider the regime where the infidelity is asymptotically small ($p\to 0$), allowing us to retain only the first-order term in $p$ while neglecting higher-order contributions. Since the protocol improves fidelity at each round, this assumption holds throughout. Achieving any finite infidelity requires only a finite number of distillation rounds, ensuring constant overhead for the yield rate.
Applying the $\left[n,n-1,2\right]$ protocol to $n$ Bell states with initial infidelity $p_{\rm in}$, the output infidelity is given by $p_{\rm{out}}=\frac{n(n-1)}{2}p_{\rm{in}}^{2}$ and the distillation success probability is $P_{S}=1-np_{\rm{in}}$, both calculated using a first-order approximation. The protocol requires $p_{\rm{out}}<p_{\rm{in}}$, which imposes the condition $n\leq\sqrt{2}p_{\rm{in}}^{-1/2}$ on the block size. To satisfy this condition, we set $n=\sqrt{2}p_{\rm{in}}^{-\alpha}$, where $\alpha$ is a constant parameter chosen within the range $0\leq\alpha\leq\frac{1}{2}$.
In a sequence of distillations, each iteration $k$ is characterized by input infidelity $p_{k-1}$ and output infidelity $p_{k}$, with the output of iteration $k$ serving as the input for iteration $k+1$. The initial infidelity is $p_{0}$, used as input to the first iteration. The recursive parameters for iteration $k$ are as follows:
$n_{k}$ $=$ $\sqrt{2}p_{k-1}^{-\alpha}\;,$
$r_{k}$ $=$ $\frac{n_{k}-1}{n_{k}}\;,$
$p_{k}$ $=$ $\frac{1}{2}n_{k}^{2}p_{k-1}^{2}\;,$
$P_{k}$ $=$ $1-n_{k}p_{k-1}\;,$
where $n_{k}$, $r_{k}$, $p_{k}$, and $P_{k}$ represent the block size, encoding rate, output infidelity, and success probability, respectively. Solving these recursive equations yields:
$n_{k}$ $=$ $\sqrt{2}p_{0}^{-\alpha(2-2\alpha)^{k-1}}\;,$
$r_{k}$ $=$ $1-\frac{\sqrt{2}}{2}p_{0}^{\alpha(2-2\alpha)^{k-1}}\;,$
$p_{k}$ $=$ $p_{0}^{(2-2\alpha)^{k}}\;,$
$P_{k}$ $=$ $1-\sqrt{2}p_{0}^{\frac{1}{2}(2-2\alpha)^{k}}\;.$
The overall yield rate is given by $R=\prod_{k}r_{k}P_{k}$. Taking its logarithm and applying the approximation $\ln(1+x)\approx x$ for small $x$, we derive the first-order result
$-\ln{R}=\sum_{k}\left(\sqrt{2}p_{0}^{\frac{1}{2}(2-2\alpha)^{k}}+\frac{\sqrt{2}}{2}p_{0}^{\alpha(2-2\alpha)^{k-1}}\right)\;.$
For $0<\alpha<\frac{1}{2}$ at large $k$, $(2-2\alpha)^{k}>k$. We can bound the summation as
$c$ $=$ $\sum_{k}\left(\sqrt{2}p_{0}^{\frac{1}{2}(2-2\alpha)^{k}}+\frac{\sqrt{2}}{2}p_{0}^{\alpha(2-2\alpha)^{k-1}}\right)$
$<$ $\sum_{k}\left(\sqrt{2}p_{0}^{\frac{1}{2}k}+\frac{\sqrt{2}}{2}p_{0}^{\alpha(k-1)}\right)$
$=$ $\frac{\sqrt{2p_{0}}}{1-\sqrt{p_{0}}}+\frac{\sqrt{2}}{2(1-p_{0}^{\alpha})}\;.$
Thus, the overall yield rate of the adaptive entanglement distillation is lower bounded by $R>e^{-c}$, ensuring a constant yield rate. Further optimization of the protocol is possible; for example, selecting block sizes as $n_{k}\sim k^{2}$ also satisfies the threshold requirements and maintains a constant lower bound for the overall yield rate.
Appendix C Integrating Stabilizer Measurements, Error Correction, and Decoding
@@
We can integrate stabilizer measurements with single-qubit measurement decoding. Consider a stabilizer code in its standard form. The $i$th generator, for $i<r$, consists of a single bit flip on the $i$th qubit. After measuring $g_{i}$, the following stabilizer and single-qubit measurements use only Pauli-$Z$ operators on the $i$th qubit. This allows us to replace $Z_{i}$ in these stabilizers with its measurement outcome, reducing the number of two-qubit gates. Refer to Figure 4 for an example of entanglement distillation using the $\left[\!\left[5,1,3\right]\!\right]$ code.
Appendix D Decoding Quantum Convolutional Codes with Laurent Series
@@
$X_{L}^{(i)}$ $=$ $\bigotimes_{l:q_{l}=1}^{m}X_{\left(l+1\right)n-k+i}\;,$
$Z_{L}^{(i)}$ $=$ $\bigotimes_{l:q_{l}=1}^{m}Z_{\left(l+1\right)n-k+i}\;.$
These operators, associated with a catastrophic encoding circuit [99], require further decoding. For example, with $Q\left(D\right)=1+D+D^{2}$, the logical qubits are encoded by the following circuit
@@
$\begin{matrix}\left(-1\right)^{a_{x}^{(1)}+a_{x}^{(2)}+b_{x}^{(1)}+b_{x}^{(2)}}X_{3}\otimes X_{3}\;,&\left(-1\right)^{a_{z}^{(1)}+a_{z}^{(2)}+b_{z}^{(1)}+b_{z}^{(2)}}Z_{3}\otimes Z_{3}\\ \left(-1\right)^{a_{x}^{(2)}+b_{x}^{(2)}}X_{3}X_{4}\otimes X_{3}X_{4}\;,&\left(-1\right)^{a_{z}^{(2)}+b_{z}^{(2)}}Z_{3}Z_{4}\otimes Z_{3}Z_{4}\\ X_{3}X_{4}X_{5}\otimes X_{3}X_{4}X_{5}\;,&Z_{3}Z_{4}Z_{5}\otimes Z_{3}Z_{4}Z_{5}\\ \vdots&\vdots\end{matrix}\;.$
Here, $a_{x}^{(1)}$ and $b_{x}^{(1)}$ are the measurement outcomes of $X_{-1}X_{1}$ by Alice and Bob, respectively. $a_{x}^{(2)}$ and $b_{x}^{(2)}$ are the outcomes of $X_{0}X_{2}$, $a_{z}^{(1)}$ and $b_{z}^{(1)}$ are the outcomes of $Z_{-1}Z_{1}$, and $a_{z}^{(2)}$ and $b_{z}^{(2)}$ are the outcomes of $Z_{0}Z_{2}$. By recursively calculating the parities for subsequent physical Bell states, these logical Bell states are decoded.
Errors in parity measurements can propagate indefinitely. For example, if an error occurs on $a_{x}^{1}$, it introduces an incorrect recovery phase for $X_{3}\otimes X_{3}$, which then propagates through subsequent operators such as $X_{4}\otimes X_{4}$ and $X_{5}\otimes X_{5}$, leading to catastrophic errors. To address this, we segment the sequence into blocks of $N$ frames. Measuring parities independently for each block confines error propagation to a length of $N$. For a general polynomial $Q\left(D\right)=\sum_{l=0}^{m}q_{l}D^{l}$, the entanglement-assisted approach consumes $m$ pre-shared Bell states and the first $m$ Bell states per block, yielding $N-m$ decoded physical Bell states. The overhead, defined as the ratio of consumed to output Bell states, is $\frac{2m}{N-m}$, approaching zero as $N$ increases.
Appendix E Deriving the Standard Form for the Rate-1/3 Convolutional Code
@@
This process results in the standard form as specified in Equation 17. Note that during Gaussian elimination, multiplication factors representing finite shifts are preferred. Using a Laurent series as a multiplication factor, such as $\frac{1}{1+D}=1+D+D^{2}+\cdots$, can result in a standard form consisting of Laurent series entries, increasing decoding complexity. Obtaining a standard form solely through row Gaussian elimination is challenging, as it requires column entries to be coprime; otherwise, only the greatest common divisor of the polynomials is obtained. Further elimination would require column operations involving two-qubit gates, which we aim to avoid in measurement-based decoding. We leave it as an open problem to develop a general algorithm for achieving a standard form with finite entries.
Appendix F Modeling Errors in Measurement-Based Decoding
@@
$X_{L}^{(i)}$ $=$ $\left(-1\right)^{\alpha_{x}^{(i)}}X_{n-k+i}\;,$
$Z_{L}^{(i)}$ $=$ $\left(-1\right)^{\alpha_{z}^{(i)}}Z_{n-k+i}\;,$
where $\alpha_{x}^{(i)}=\sum_{j:C_{ij}^{T}=1}a_{j}+\sum_{j:E_{ij}^{T}=1}a_{r+j}$ and $\alpha_{z}^{(i)}=\sum_{j:\left(A_{2}^{T}\right)_{ij}=1}a_{j}$, as defined in Equations 10 and 11. Noise from single-qubit measurements leads to phase errors. $\alpha_{x}^{(i)}$ errors yield a $Z$ error, $\alpha_{z}^{(i)}$ errors yield an $X$ error, and errors in both $\alpha_{x}^{(i)}$ and $\alpha_{z}^{(i)}$ yield a $Y$ error on the decoded qubit.